# Picky elements, subnormalisers, and character correspondences

### v1

@@ 3.3. Non-simply laced and twisted types
There is nothing to show for $m=1$, so assume $m\geq 2$. Let $x\in G$ be unipotent and denote by $\lambda_{1}\geq\lambda_{2}\geq\cdots$ the lengths of its Jordan blocks. According to the normal forms given in [7, 4.1.3], $x$ has a conjugate in $B$ in which the entry at position $(i,i+1)$ is zero for at least $2m-\lambda_{1}$ indices $1\leq i\leq m$. Assume $12m-\lambda_{1}>$ 1. Then the image of $x$ in the standard Levi subgroup ${\operatorname{GL}}_{m}(q^{2})$ is not regular and thus ${\operatorname{Sub}}_{G}(x)$ contains the corresponding maximal parabolic subgroup by Propositions 3.5 and 3.6. Furthermore, the image of  $x$ in the standard Levi subgroup ${\operatorname{GU}}_{2m-2}(q)$ also has at least 2 zero entries directly above the main diagonal, unless $\lambda_{1}=\lambda_{2}$. In the latter case we must have $42m=$ 4, and then the image of $x$ in ${\operatorname{GU}}_{2m-2}(q)={\operatorname{GU}}_{2}(q)$ is trivial. So in either case, by induction ${\operatorname{Sub}}_{G}(x)$ also contains this end node parabolic subgroup and thus equals $G$.
This only leaves the cases when $2m-\lambda_{1}\leq 1$, that is, $x$ has Jordan type $(2m)$ or $(2m-1,1)$. The first of these is the class of regular unipotent elements. If $x\in U$ has a Jordan block of length $12m-$ 1, then it must involve elements from all simple root subgroups except possibly for the ’middle’ one. In particular, the only standard parabolic subgroups whose unipotent radical could contain $x$ are $B$ and $P_{m}$. It is easy to see that ${\mathbf{O}}_{p}(P_{m})$ contains elements with a Jordan block of length $12m-$ 1; then Lemma 2.10 shows ${\operatorname{Sub}}_{G}(x)=P_{m}$. ∎
@@ 4. On Conjecture 1 for the defining prime
Let $x\in B\leq G$ be regular unipotent. Under our assumptions on ${\mathbf{G}}$ and $p$, by the theorem of Green–Lehrer–Lusztig [4, Cor. 8.3.6] the irreducible characters of $G$ that do not vanish on $x$ are exactly those of degree prime to $p$, there are $|{\mathbf{Z}}(G)|q^{l}$ of these, where $l$ denotes the semisimple rank of ${\mathbf{G}}$, and they all take value ±$\pm 1$ on $x$. This implies also that $x$ must be rational. Now ${\operatorname{Sub}}_{G}(x)=B$ as $x$ is picky by Theorem 3.2. By the proven McKay conjecture for groups of Lie type in defining characteristic [26] there is the same number of irreducible characters of $B$ of $p^{\prime}$-degree. Moreover, as $x$ is picky, by Lemma 2.1 it must also be rational in $B$ and $|{\mathbf{C}}_{B}(x)|=|{\mathbf{C}}_{G}(x)|$. This implies that the $|{\mathbf{Z}}(G)|q^{l}$ irreducible characters of $B$ not vanishing on $x$ must also take values ±$\pm 1$. ∎
@@
The character tables for $G=G_{2}(2^{f})$ and a Borel subgroup $B$ of $G$ were determined by Enomoto–Yamada [11]. First let $x\in B$ be regular (there are two such classes) so ${\operatorname{Sub}}_{G}(x)=B$. According to Tables I and IV of that paper both ${\operatorname{Irr}}^{x}(G)$ and ${\operatorname{Irr}}^{x}(B)$ consist of $q^{2}$ characters of $p^{\prime}$-degree and 4 characters with $p$-part of the degree equal to $q/2$. The values of the former on $x$ are ±$\pm 1$, and ±$\pm q/2$ for the latter in both groups, and their rationality properties agree (for the $p^{\prime}$-characters this follows by [28]), so Conjecture 1+ holds for $x$.
The character table of $G=G_{2}(3^{f})$ and of its Borel subgroup can be found in [10, Tab. I and VII]. The group $G$ contains three classes of regular unipotent elements $x$, where again ${\operatorname{Sub}}_{G}(x)\sim_{G}B$. Here, both ${\operatorname{Irr}}^{x}(G)$ and ${\operatorname{Irr}}^{x}(B)$ consist of $q^{2}$ characters of $p^{\prime}$-degree and 6 characters with $p$-part of the degree equal to $q/3$. The values of the $p^{\prime}$-characters on $x$ are all ±$\pm 1$. The other characters take values ±$\pm 2q/3$ (4 times) and ±$\pm q/3$ (twice) on one of the classes, on the other two the values are ±$\pm q/3$ (4 times) and $q/3+\zeta_{3}^{i}q$ for $i=1,2$, where $\zeta$ is a primitive third root of unity, both for $G$ and for $B$. Again, Conjecture 1+ is seen to hold.
@@
$\begin{array}[]{c|cccccccc}&\theta_{1},\theta_{1}^{\prime}&\theta_{2},\theta_{2}^{\prime}&\theta_{3}&\theta_{4},\theta_{9}(\pm 1)&\theta_{8}&\chi_{2}(k)&\chi_{2}^{\prime}(k)\\ \#&2&2&1&3&1&{\frac{1}{2}}(q{\!-\!}3{\!-\!}\epsilon)&{\frac{1}{2}}(q{\!-\!}1{\!+\!}\epsilon)\\ \hline\cr A_{4}&\frac{1}{6}q(\epsilon q-1)&-{\frac{1}{2}}q(\epsilon q-1)&\frac{1}{3}q(\epsilon q-1)&\frac{1}{3}q(\epsilon q+2)&q&q&-q\\ \end{array}$
@@
$\begin{array}[]{c|ccccc}&\theta_{3}(\pm 1)&\theta_{2}(\pm 1)&\theta_{4}&\theta_{5},\theta_{6}(\pm 1)&\chi_{3}(k)\\ \#&2&2&1&3&q{\!-\!}1\\ \hline\cr A_{5}&\frac{1}{6}q(\epsilon q-1)&-{\frac{1}{2}}q(\epsilon q-1)&\frac{1}{3}q(\epsilon q-1)&\frac{1}{3}q(\epsilon q+2)&q\\ \end{array}$
@@
For $p=2$, the character table of $G$ was computed in [8] and the one of a Borel subgroup in [16, Tab. A.6]. For both groups there are $q^{4}$ characters of odd degree, all taking value ±$\pm 1$ on regular unipotent elements, and four characters of degree divisible by $q^{3}$, taking values ±$\pm q^{2}/2$.
Now we consider the elements $x$ in the subregular unipotent class $D_{4}(a_{1})$ of $G$ with ${\operatorname{Sub}}_{G}(x)=P_{2}$, the second maximal standard parabolic subgroup. Here, by the tables in [8, 15, 16] both $G$ and $P_{2}$ possess $q^{3}$ irreducible characters of degree divisible by $p$ (in fact, precisely by $q$) not vanishing on $x$, and all of them take value ±$\pm q$ on $x$. Conjecture 1+ now follows, using [28] for the character fields of $p^{\prime}$-characters. ∎
@@
The character table of $G:={}^{2}G_{2}(q^{2})$ was found by Ward [36] while that of a Sylow 3-normaliser was computed by van der Waall [35, p. 173]. In Table 3 we reproduce the values of those characters in ${\operatorname{Irr}}(G)$ respectively ${\operatorname{Irr}}(B)$ of degree divisible by $p$ not vanishing on some non-trivial unipotent element.
@@
$\begin{array}[]{c|cccc||c|cccc}{}^{2}G_{2}(q^{2})&1&X&T&YT^{i}&\qquad B&1&X&T&YT^{i}\\ \hline\cr\overline{\xi_{5}},\xi_{7}&{\frac{1}{2}}{\bar{q}}\Phi_{1}\Phi_{6}^{\prime\prime}&-{\frac{1}{2}}(q^{2}+{\bar{q}})&{\bar{q}}b&{\bar{q}}\zeta_{3}^{i}&\psi_{3},\overline{\psi_{5}}&{\frac{1}{2}}{\bar{q}}\Phi_{1}&{\frac{1}{2}}{\bar{q}}\Phi_{1}&{\bar{q}}b&{\bar{q}}\zeta_{3}^{i}\\ \overline{\xi_{6}},\xi_{8}&{\frac{1}{2}}{\bar{q}}\Phi_{1}\Phi_{6}^{\prime}&{\frac{1}{2}}(q^{2}-{\bar{q}})&{\bar{q}}b&{\bar{q}}\zeta_{3}^{i}&\psi_{4},\overline{\psi_{6}}&{\frac{1}{2}}{\bar{q}}\Phi_{1}&{\frac{1}{2}}{\bar{q}}\Phi_{1}&{\bar{q}}b&{\bar{q}}\zeta_{3}^{i}\\ \overline{\xi_{9}},\xi_{10}&{\bar{q}}(q^{4}-1)&-{\bar{q}}&2{\bar{q}}b&-{\bar{q}}\zeta_{3}^{i}&\psi_{1},\overline{\psi_{2}}&{\bar{q}}\Phi_{1}&{\bar{q}}\Phi_{1}&2{\bar{q}}b&-{\bar{q}}\zeta_{3}^{i}\\ \xi_{4}&q^{2}\Phi_{6}&q^{2}&.&.&\psi&q^{2}\Phi_{1}&-q^{2}&.&.\\ \end{array}$
@@
We need to consider the regular classes, the picky subregular classes from Theorem 3.2(5) as well as the classes with proper subnormaliser in Proposition 3.7(5). The values of unipotent characters of $G$ are given in [21], the values of all other characters at least on unipotent elements can be found in [13]. The character table of a Borel subgroup of $G$ was computed in [17]. For $x$ in one of the four regular unipotent classes or the two subregular picky classes, ${\operatorname{Irr}}^{x}(G)$ and ${\operatorname{Irr}}^{x}(B)$ both consists of $q^{4}$ characters of odd degree, $2q^{2}$ characters whose degree has 2-part $\sqrt{2}q/2$, and 8 characters whose degree has 2-part $q^{4}/4$. These take values ±$\pm 1$, $q\sqrt{-2}/2$, and ±$\pm q^{2}/2$, ±$\pm\sqrt{-1}q^{2}/2$ on regular elements, and values ±$\pm 1$, $q\sqrt{2}/2$ respectively ±$\pm\sqrt{2}q^{3}/4$ on the subregular elements, both for $G$ and $B$. A bijection preserving character fields over ${\mathbb{Q}}_{p}$ was given in [20].
@@
Table 4. $2$-Parts of character values for ${}^{2}\!F_{4}(q^{2})$
#$\begin{array}[]{c|c|ccccccccc}&\#&1&u_{9}&u_{10}&u_{11}=u_{12}^{-1}&u_{13-14}&u_{15-18}\\ \hline\cr\chi_{2,3,23,24}&2q^{2}&{\bar{q}}&{\bar{q}}&{\bar{q}}&{\bar{q}}&{\bar{q}}&{\bar{q}}\\ \chi_{4,27,30,33}&q^{2}&q^{2}&q^{2}&q^{2}&q^{2}&.&.\\ \chi_{5,6,8,9}&4&{\bar{q}}^{4}&{\bar{q}}^{4}&{\bar{q}}^{4}&{\bar{q}}^{4}&{\bar{q}}^{3}&{\bar{q}}^{2}\\ \chi_{11-14}&4&{\bar{q}}^{4}&{\bar{q}}^{4}&{\bar{q}}^{4}&({\bar{q}}^{4}\pm 2i{\bar{q}}^{3})_{2}&{\bar{q}}^{3}&{\bar{q}}^{2}\\ \chi_{7,10}&2&2{\bar{q}}^{4}&2{\bar{q}}^{4}&2{\bar{q}}^{4}&2{\bar{q}}^{4}&.&.\\ \chi_{15-17}&3&q^{4}&2q^{4}&.&.&.&.\\ \end{array}$
@@
Here, Proposition 4.1 does not apply as for odd $p$ the centre of the algebraic group ${\mathbf{G}}={\operatorname{Sp}}_{4}$ is disconnected, and the prime $p=2$ is bad for ${\mathbf{G}}$. For $p=2$ the character tables of $G$ and $B$ were calculated by Enomoto [9, Tab. I and IV]. There are $q^{2}$ irreducible characters of odd degree, taking value ±$\pm 1$ on a regular unipotent element $x$, and 4 characters of degree divisible by $q/2$ taking values ±$\pm q/2$, both for $G$ and for $B$.
The character table of $G={\operatorname{Sp}}_{4}(q)$ for odd $q$ was determined by Srinivasan [34] and for $B$ by Yamada [37, Tab. I-2]. The characters not vanishing on a regular unipotent element $x\in G$ for both $G$ and $B$ are all of $p^{\prime}$-degree, and $q(q-1)$ of them take values ±$\pm 1$ on $x$, the other $4q$ take values ±$\pm(1\pm\sqrt{q^{*}})/2$, where $q^{*}:=(-1)^{(q-1)/2}q$.
Finally, for $p$ odd and $x$ the subregular element as in Proposition 3.7(2) we have ${\operatorname{Sub}}_{G}(x)=P_{1}$. The character table of $P_{1}$ can be found in [37, II-2], where the corresponding class is denoted $A_{3}(1)$. The characters of degree divisible by $p$ take values ±$\pm q$ ($q+3$ times) and ±$\pm 2q$ ($(q-1)/2$ times) on $x$, for both $G$ and $P_{1}$. ∎
@@ 5. Picky semisimple elements in groups of Lie type
We now turn to the investigations of semisimple elements of groups of Lie type, that is to say, elements of prime power order different from the defining characteristic. We stay in the setting from the beginning of Section 3, with ${\mathbf{G}}$ connected reductive and $G={\mathbf{G}}^{F}$ a finite groups of Lie type, and let $\ell\neq p$ be a prime. To investigate the existence of picky $\ell$-elements in simple groups $S$, by Lemma 2.4, we may consider ${\mathbf{G}}$ of simply connected type such that $G/{\mathbf{Z}}(G)\cong S$.
5.1. Semisimple picky elements
@@
In the above setting, either $\ell$ divides a unique cyclotomic factor $\Phi_{d}(q)$ occurring in the order polynomial of $G={\mathbf{G}}^{F}$, or Sylow $\ell$-subgroups of $G$ are non-abelian.
@@
If $\ell$ divides at least two cyclotomic factors, then it divides the order of the Weyl group of ${\mathbf{G}}$. But in this case, by a straightforward inspection of the Dynkin diagrams of simple groups, ${\mathbf{G}}$ has a semisimple subgroup of type $A_{\ell-1}$. But the groups isogenous to ${\operatorname{SL}}_{\ell}(\pm q)$ are easily seen to have non-abelian Sylow $\ell$-subgroups. ∎
Note that we do not claim that the simple groups $S=G/{\mathbf{Z}}(G)$ satisfy this dichotomy; counter-examples occur for ${\operatorname{SL}}_{2}(q)$ at $\ell=2$ and ${\operatorname{SL}}_{3}(\pm q)$ at $\ell=3$.
@@
In the above setting, assume that $\ell$ divides a unique cyclotomic factor $\Phi_{d}(q)$ in the order polynomial of $G={\mathbf{G}}^{F}$. Then an $\ell$-element $x\in G$ is picky if and only if ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}})$ where ${\mathbf{S}}\leq{\mathbf{G}}$ is a Sylow $d$-torus with $x\in{\mathbf{S}}^{F}$.
@@
If $\ell$ divides a unique cyclotomic factor $\Phi_{d}(q)$ in the order polynomial of $G$ then any Sylow $\ell$-subgroup $P$ of $G$ must lie in a Sylow $d$-torus ${\mathbf{S}}$ of ${\mathbf{G}}$. Conversely, if ${\mathbf{S}}$ is a Sylow $d$-torus containing $P$, then ${\mathbf{S}}\leq{\mathbf{C}}_{\mathbf{G}}(P)$, hence even ${\mathbf{S}}\leq{\mathbf{H}}:={\mathbf{C}}_{\mathbf{G}}^{\circ}(P)$. Also, by inspection of the order formulae [24, Tab. 24.1] our condition on $\ell$ implies that $\ell$ does not divide $|W|$. Thus, $\ell$ is not a torsion prime for ${\mathbf{G}}$ and then neither for ${\mathbf{H}}$, so $P\leq{\mathbf{Z}}^{\circ}({\mathbf{H}})=:{\mathbf{S}}^{\prime}$, a torus. Since $\ell$ divides a unique $\Phi_{d}(q)$ this implies that ${\mathbf{S}}\leq{\mathbf{S}}^{\prime}$, i.e., ${\mathbf{S}}$ is the Sylow $d$-torus of ${\mathbf{Z}}^{\circ}({\mathbf{C}}_{\mathbf{G}}^{\circ}(P))$ and thus uniquely determined by $P$.
Hence $x\in G$ lies in a unique Sylow $\ell$-subgroup of $G$ if and only if it lies in a unique Sylow $d$-torus ${\mathbf{S}}$ of ${\mathbf{G}}$, if and only if ${\mathbf{S}}$ is the unique Sylow $d$-torus of ${\mathbf{C}}_{\mathbf{G}}(x)$, if and only if ${\mathbf{S}}\leq{\mathbf{Z}}({\mathbf{C}}_{\mathbf{G}}(x))$, if and only if ${\mathbf{C}}_{\mathbf{G}}(x)\leq{\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$. As $x\in{\mathbf{S}}$ we also have ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})\leq{\mathbf{C}}_{\mathbf{G}}(x)$, so ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})={\mathbf{C}}_{\mathbf{G}}(x)$. This implies of course ${\mathbf{C}}_{G}({\mathbf{S}})={\mathbf{C}}_{G}(x)$. Conversely, assume ${\mathbf{C}}_{G}({\mathbf{S}})={\mathbf{C}}_{G}(x)$ but ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})<{\mathbf{C}}_{\mathbf{G}}(x)$. Since ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$ is a Levi subgroup, hence connected, this means that either ${\mathbf{C}}_{\mathbf{G}}(x)$ is disconnected, or of strictly larger dimension than ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$. The first is not possible as $\ell$ is not a torsion prime. In the second case, as ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$ contains a maximal torus of ${\mathbf{G}}$, the maximal unipotent subgroups of ${\mathbf{C}}_{\mathbf{G}}(x)$ must have larger dimension than those of ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$. But then $|{\mathbf{C}}_{G}(x)|_{p}>|{\mathbf{C}}_{G}({\mathbf{S}})|_{p}$, a contradiction. In conclusion, $x\in G$ is picky if and only if ${\mathbf{C}}_{G}({\mathbf{S}})={\mathbf{C}}_{G}(x)$, for ${\mathbf{S}}$ the (unique) Sylow $d$-torus containing $x$. ∎

### v2

@@ 3.3. Non-simply laced and twisted types
There is nothing to show for $m=1$, so assume $m\geq 2$. Let $x\in G$ be unipotent and denote by $\lambda_{1}\geq\lambda_{2}\geq\cdots$ the lengths of its Jordan blocks. According to the normal forms given in [7, 4.1.3], $x$ has a conjugate in $B$ in which the entry at position $(i,i+1)$ is zero for at least $2m-\lambda_{1}$ indices $1\leq i\leq m$. Assume $2m-\lambda_{1}>1$. Then the image of $x$ in the standard Levi subgroup ${\operatorname{GL}}_{m}(q^{2})$ is not regular and thus ${\operatorname{Sub}}_{G}(x)$ contains the corresponding maximal parabolic subgroup by Propositions 3.5 and 3.6. Furthermore, the image of  $x$ in the standard Levi subgroup ${\operatorname{GU}}_{2m-2}(q)$ also has at least 2 zero entries directly above the main diagonal, unless $\lambda_{1}=\lambda_{2}$. In the latter case we must have $2m=4$, and then the image of $x$ in ${\operatorname{GU}}_{2m-2}(q)={\operatorname{GU}}_{2}(q)$ is trivial. So in either case, by induction ${\operatorname{Sub}}_{G}(x)$ also contains this end node parabolic subgroup and thus equals $G$.
This only leaves the cases when $2m-\lambda_{1}\leq 1$, that is, $x$ has Jordan type $(2m)$ or (2m1,1)(2m-1,1). The first of these is the class of regular unipotent elements. If $x\in U$ has a Jordan block of length 2m12m-1, then it must involve elements from all simple root subgroups except possibly for the ’middle’ one. In particular, the only standard parabolic subgroups whose unipotent radical could contain $x$ are $B$ and $P_{m}$. It is easy to see that ${\mathbf{O}}_{p}(P_{m})$ contains elements with a Jordan block of length 2m12m-1; then Lemma 2.14 shows ${\operatorname{Sub}}_{G}(x)=P_{m}$. ∎
@@ 4. On Conjecture 1 for the defining prime
Let $x\in B\leq G$ be regular unipotent. Under our assumptions on $\mathbf{G}$} and $p$, by the theorem of Green–Lehrer–Lusztig [4, Cor. 8.3.6] the irreducible characters of $G$ that do not vanish on $x$ are exactly those of degree prime to $p$, there are $|{\mathbf{Z}}(G)|q^{l}$ of these, where $l$ denotes the semisimple rank of $\mathbf{G}$}, and they all take value $\pm 1$ on $x$. This implies also that $x$ must be rational. Now ${\operatorname{Sub}}_{G}(x)=B$ as $x$ is picky by Theorem 3.2. By the proven McKay conjecture for groups of Lie type in defining characteristic [29] there is the same number of irreducible characters of $B$ of $p^{\prime}$-degree. Moreover, as $x$ is picky, by Lemma 2.1 it must also be rational in $B$ and $|{\mathbf{C}}_{B}(x)|=|{\mathbf{C}}_{G}(x)|$. This implies that the $|{\mathbf{Z}}(G)|q^{l}$ irreducible characters of $B$ not vanishing on $x$ must also take values $\pm 1$. ∎
@@
The character tables for $G=G_{2}(2^{f})$ and a Borel subgroup $B$ of $G$ were determined by Enomoto–Yamada [11]. First let $x\in B$ be regular (there are two such classes) so ${\operatorname{Sub}}_{G}(x)=B$. According to Tables I and IV of that paper both ${\operatorname{Irr}}^{x}(G)$ and ${\operatorname{Irr}}^{x}(B)$ consist of $q^{2}$ characters of $p^{\prime}$-degree and 4 characters with $p$-part of the degree equal to $q/2$. The values of the former on $x$ are $\pm 1$, and $\pm q/2$ for the latter in both groups, and their rationality properties agree (for the $p^{\prime}$-characters this follows by [31]), so Conjecture 1+ holds for $x$.
The character table of $G=G_{2}(3^{f})$ and of its Borel subgroup can be found in [10, Tab. I and VII]. The group $G$ contains three classes of regular unipotent elements $x$, where again ${\operatorname{Sub}}_{G}(x)\sim_{G}B$. Here, both ${\operatorname{Irr}}^{x}(G)$ and ${\operatorname{Irr}}^{x}(B)$ consist of $q^{2}$ characters of $p^{\prime}$-degree and 6 characters with $p$-part of the degree equal to $q/3$. The values of the $p^{\prime}$-characters on $x$ are all $\pm 1$. The other characters take values $\pm 2q/3$ (4 times) and $\pm q/3$ (twice) on one of the classes, on the other two the values are $\pm q/3$ (4 times) and $q/3+\zeta_{3}^{i}q$ for $i=1,2$, where $\zeta$ is a primitive third root of unity, both for $G$ and for $B$. Again, Conjecture 1+ is seen to hold.
@@
[…]
@@
[…]
@@
For $p=2$, the character table of $G$ was computed in [8] and the one of a Borel subgroup in [16, Tab. A.6]. For both groups there are $q^{4}$ characters of odd degree, all taking value $\pm 1$ on regular unipotent elements, and four characters of degree divisible by $q^{3}$, taking values $\pm q^{2}/2$.
Now we consider the elements $x$ in the subregular unipotent class $D_{4}(a_{1})$ of $G$ with ${\operatorname{Sub}}_{G}(x)=P_{2}$, the second maximal standard parabolic subgroup. Here, by the tables in [8, 15, 16] both $G$ and $P_{2}$ possess $q^{3}$ irreducible characters of degree divisible by $p$ (in fact, precisely by $q$) not vanishing on $x$, and all of them take value $\pm q$ on $x$. Conjecture 1+ now follows, using [31] for the character fields of $p^{\prime}$-characters. ∎
@@
The character table of $G:={}^{2}G_{2}(q^{2})$ was found by Ward [39] while that of a Sylow 3-normaliser was computed by van der Waall [38, p. 173]. In Table 3 we reproduce the values of those characters in ${\operatorname{Irr}}(G)$ respectively ${\operatorname{Irr}}(B)$ of degree divisible by $p$ not vanishing on some non-trivial unipotent element.
@@
[…]
@@
We need to consider the regular classes, the picky subregular classes from Theorem 3.2(5) as well as the classes with proper subnormaliser in Proposition 3.7(5). The values of unipotent characters of $G$ are given in [22], the values of all other characters at least on unipotent elements can be found in [13]. The character table of a Borel subgroup of $G$ was computed in [17]. For $x$ in one of the four regular unipotent classes or the two subregular picky classes, ${\operatorname{Irr}}^{x}(G)$ and ${\operatorname{Irr}}^{x}(B)$ both consists of $q^{4}$ characters of odd degree, $2q^{2}$ characters whose degree has 2-part $\sqrt{2}q/2$, and 8 characters whose degree has 2-part $q^{4}/4$. These take values $\pm 1$, $q\sqrt{-2}/2$, and $\pm q^{2}/2$, $\pm\sqrt{-1}q^{2}/2$ on regular elements, and values $\pm 1$, $q\sqrt{2}/2$ respectively $\pm\sqrt{2}q^{3}/4$ on the subregular elements, both for $G$ and $B$. A bijection preserving character fields over ${\mathbb{Q}}_{p}$ was given in [20].
@@
Table 4. 22-Parts of character values for $^{2}\!F_{4}(q^{2})$
[…]
@@
Here, Proposition 4.1 does not apply as for odd $p$ the centre of the algebraic group ${\mathbf{G}}={\operatorname{Sp}}_{4}$ is disconnected, and the prime $p=2$ is bad for $\mathbf{G}$}. For $p=2$ the character tables of $G$ and $B$ were calculated by Enomoto [9, Tab. I and IV]. There are $q^{2}$ irreducible characters of odd degree, taking value $\pm 1$ on a regular unipotent element $x$, and 4 characters of degree divisible by $q/2$ taking values $\pm q/2$, both for $G$ and for $B$.
The character table of $G={\operatorname{Sp}}_{4}(q)$ for odd $q$ was determined by Srinivasan [37] and for $B$ by Yamada [40, Tab. I-2]. The characters not vanishing on a regular unipotent element $x\in G$ for both $G$ and $B$ are all of $p^{\prime}$-degree, and q(q1)q(q-1) of them take values $\pm 1$ on $x$, the other $4q$ take values $\pm(1\pm\sqrt{q^{*}})/2$, where $q^{*}:=(-1)^{(q-1)/2}q$.
Finally, for $p$ odd and $x$ the subregular element as in Proposition 3.7(2) we have ${\operatorname{Sub}}_{G}(x)=P_{1}$. The character table of $P_{1}$ can be found in [40, II-2], where the corresponding class is denoted $A_{3}(1)$. The characters of degree divisible by $p$ take values $\pm q$ ($q+3$ times) and $\pm 2q$ ((q1)/2(q-1)/2 times) on $x$, for both $G$ and $P_{1}$. ∎
@@ 5. Picky semisimple elements in groups of Lie type
We now turn to the investigation of semisimple elements of groups of Lie type, that is to say, elements of (prime power) order different from the defining characteristic. We stay in the setting from the beginning of Section 3, with $\mathbf{G}$} connected reductive and $G={\mathbf{G}}^{F}$ a finite groups of Lie type, and let $\ell\neq p$ be a prime. To investigate the existence of picky $\ell$-elements in simple groups $S$, by Lemma 2.4, we may consider $\mathbf{G}$} of simply connected type such that $G/{\mathbf{Z}}(G)\cong S$.
5.1. The case of abelian Sylow $\ell$-subgroups
@@
In the above setting, if $\ell$ divides $\Phi_{d}(q)$ for a unique cyclotomic factor $\Phi_{d}$ occurring in the order polynomial of $({\mathbf{G}},F)$, then Sylow $\ell$-subgroups of $G$ are abelian. If $\mathbf{G}$} is simple, the converse holds.
This is shown in [26, Prop. 2.2]. Note that we do not claim that the simple groups $S=G/{\mathbf{Z}}(G)$ satisfy this dichotomy; counter-examples occur for ${\operatorname{SL}}_{2}(q)$ at $\ell=2$ and ${\operatorname{SL}}_{3}(\pm q)$ at $\ell=3$. The following is an analogue of Proposition 2.6:
Proposition 5.2.
In the above setting, assume that $\ell$ divides a unique cyclotomic factor $\Phi_{d}(q)$ in the order polynomial of $({\mathbf{G}},F)$. Let $x\in G$ be an $\ell$-element. Then ${\operatorname{Sub}}_{G}(x)$ is generated by the $F$-fixed points of the normalisers of the Sylow $d$-tori of $\mathbf{G}$} containing $x$.
@@
Let $x\in G$ be an $\ell$-element, and $P$ a Sylow $\ell$-subgroup of $G$ containing $x$. Then $P$ is abelian by Lemma 5.1 and lies in a Sylow $d$-torus $\mathbf{S}$} of $\mathbf{G}$}. Conversely, if $\mathbf{S}$} is a Sylow $d$-torus containing $P$, then ${\mathbf{S}}\leq{\mathbf{C}}_{\mathbf{G}}(P)$, hence even ${\mathbf{S}}\leq{\mathbf{H}}:={\mathbf{C}}_{\mathbf{G}}^{\circ}(P)$. By inspection of the order formulae [27, Tab. 24.1] our condition on $\ell$ implies that $\ell$ does not divide $|W|$. Thus, $\ell$ is not a torsion prime for $\mathbf{G}$} and then neither for $\mathbf{H}$}, so $P\leq{\mathbf{Z}}^{\circ}({\mathbf{H}})=:{\mathbf{S}}^{\prime}$, a torus. Since $\ell$ divides a unique $\Phi_{d}(q)$ this implies that ${\mathbf{S}}\leq{\mathbf{S}}^{\prime}$, i.e., $\mathbf{S}$} is the Sylow $d$-torus of ${\mathbf{Z}}^{\circ}({\mathbf{C}}_{\mathbf{G}}^{\circ}(P))$ and thus uniquely determined by $P$. Since $P$ is characteristic in ${\mathbf{S}}^{F}$ we have $N_{G}(P)=N_{G}({\mathbf{S}}^{F})$, and our claim follows by Proposition 2.6. ∎
@@
In the above setting, assume that $\ell$ divides a unique cyclotomic factor $\Phi_{d}(q)$ in the order polynomial of $({\mathbf{G}},F)$. Then an $\ell$-element $x\in G$ is picky if and only if ${\mathbf{C}}_{G}(x)={\mathbf{C}}_{G}({\mathbf{S}})$ where ${\mathbf{S}}\leq{\mathbf{G}}$ is a Sylow $d$-torus with $x\in{\mathbf{S}}^{F}$.
@@
By Lemma 5.1 we are in the situation of Proposition 5.2. Hence $x\in G$ lies in a unique Sylow $\ell$-subgroup of $G$ if and only if it lies in a unique Sylow $d$-torus $\mathbf{S}$} of $\mathbf{G}$}, if and only if $\mathbf{S}$} is the unique Sylow $d$-torus of ${\mathbf{C}}_{\mathbf{G}}(x)$, if and only if ${\mathbf{S}}\leq{\mathbf{Z}}({\mathbf{C}}_{\mathbf{G}}(x))$, if and only if ${\mathbf{C}}_{\mathbf{G}}(x)\leq{\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$. As $x\in{\mathbf{S}}$ we also have ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})\leq{\mathbf{C}}_{\mathbf{G}}(x)$, so ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})={\mathbf{C}}_{\mathbf{G}}(x)$. This implies of course ${\mathbf{C}}_{G}({\mathbf{S}})={\mathbf{C}}_{G}(x)$. Conversely, assume ${\mathbf{C}}_{G}({\mathbf{S}})={\mathbf{C}}_{G}(x)$ but ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})<{\mathbf{C}}_{\mathbf{G}}(x)$. Since ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$ is a Levi subgroup, hence connected, this means that either ${\mathbf{C}}_{\mathbf{G}}(x)$ is disconnected, or of strictly larger dimension than ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$. The first is not possible as $\ell$ is not a torsion prime. In the second case, as ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$ contains a maximal torus of $\mathbf{G}$}, the maximal unipotent subgroups of ${\mathbf{C}}_{\mathbf{G}}(x)$ must have larger dimension than those of ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$. But then $|{\mathbf{C}}_{G}(x)|_{p}>|{\mathbf{C}}_{G}({\mathbf{S}})|_{p}$, a contradiction. In conclusion, $x\in G$ is picky if and only if ${\mathbf{C}}_{G}({\mathbf{S}})={\mathbf{C}}_{G}(x)$, for $\mathbf{S}$} the (unique) Sylow $d$-torus containing $x$. ∎